OpenStudy (anonymous):

find all the real zeros of the polynomial f(x)=x^4-74x^2+1225 and determine the multiplicity of each

OpenStudy (anonymous):

Factor: (x^2 - 49)(x^2 - 25) = 0 set each factor equal to zero, you get x^2 - 49 = 0 or x = 7 and -7 x^2 - 25 = 0, you get x = -5 and 5 So 4 solutions for x: -7,-5,5,7 The catch was to be able to factor this polynomial.

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

welcome.